# Stable Symmetry-Protected Topological Phases in Systems with Heralded Noise

### v1

@@ Abstract
Figure 1: (a) Top: schematic showing stabilizers $X_{j-1}Z_{j}X_{j+1}$ (triangles) and locations of erasure errors (flags) relative to the qubits (circles). Bottom: the correction protocol for different configurations (the remainder of the system is unchanged). The left displays processes that move a stabilizer defect (filled triangle) past an adjacent erasure (filled flag); the right shows processes that shorten erasure strings which do not end on a stabilizer defect. (b) The steady-state expectation values of the string-order ($\Omega$), density of stabilizer defects ($n_{d}$), and the density of erasures ($n_{e}$) are shown as a function of the erasure noise rate $\eta/\gamma$ and clearly display a sharp transition. The data is shown for $L=128,256,512$ sites per sublattice (increasing system sizes from lighter to darker shades). The steady-state values are obtained by averaging over $5-10\times 10^{3}$ independent Monte Carlo trajectories initialized in the cluster state and simulated upto $2\times L^{1.5}$ MC sweeps.
@@
In this work, we show that for a particular class of experimentally relevant error models, strictly local error correction can stabilize one-dimensional symmetry protected topological (SPT) order in steady-state phases. These steady-state phases have a non-vanishing string-order parameter, ensuring that the mixed state remains a resource for quantum teleportation [37] and measurement-based quantum computation (MBQC) [38, 39]. In closed quantum systems, SPT order is stable against symmetric local perturbations, for which symmetry defects are locally created in pairs. In open quantum systems, only “strong” symmetries correspond to the conservation of symmetry charges within the system [40] and are hence required for defining mixed-state SPT order  [^1] Mixed-state order can be defined for $G\times H$ symmetric systems, where $G$ is a “weak” symmetry and $H$ is strong [41, 15, 42, 43, 44, 45, 46, 47, 48]. [49, 50, 51, 52, 41, 15, 42, 43, 53]. While pure SPT states are stable to strongly symmetric finite-depth local channels [49, 50], here we demonstrate that mixed SPT phases can arise as stable steady-states under local Lindbladian evolution.
@@
Background on SPTs.—We will focus on stabilizing the 1D cluster state |$\ket{\psi}_{C}=\prod_{j=1}^{2L}CX_{j,j+1}\ket{\uparrow}^{\otimes 2L}$ ($CX_{i,j}$ is a two-qubit conditional unitary), which is an SPT state protected by internal $G=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. |$\ket{\psi}_{C}$ is a representative stabilizer state within this SPT phase, satisfying $S_{j}\ket{\psi}_{C}=\ket{\psi}_{C}$ for all $j$, with stabilizers $S_{j}=X_{j-1}Z_{j}X_{j+1}$. While |$\ket{\psi}_{C}$ is short-range entangled, it cannot be connected to a $G$-symmetric paramagnetic ground state (i.e. a symmetric product state) through a $G$-symmetric finite-depth local unitary [61, 62, 63]. The resulting SPT order is encoded in the nonlocal “string-order parameters,” $\Omega^{a}_{i_{0},j_{0}}=X_{2i_{0}-1}\left(\prod_{i=i_{0}}^{j_{0}}Z_{2i}\right)X_{2j_{0}+1}$, $\Omega^{b}_{i_{0},j_{0}}=X_{2i_{0}}\left(\prod_{i=i_{0}}^{j_{0}}Z_{2i+1}\right)X_{2j_{0}+2}$, with global symmetry generators $\mathbb{Z}_{2}^{(a)}=\prod_{i=1}^{L}Z_{2i}$, $\mathbb{Z}_{2}^{(b)}=\prod_{i=1}^{L}Z_{2i-1}$ for a system with $L$ unit cells. At generic points within the phase, when the interval $(i_{0},j_{0})$ spans the entire system, and thus has a definite symmetry charge, the string operators reduce to long-range two-point correlation functions of operators localized at the boundaries, resulting in edge states with fractionalized symmetry (see Supplemental Material (SM) [64] for details). This leads to entanglement features that are a resource for teleportation or MBQC throughout the SPT phase [65, 66, 67, 68, 69, 70, 37, 71].
Noise Model.—We consider a biased erasure noise model which naturally exhibits strong $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry and which is specifically motivated by the Rydberg atom setup considered in Ref. [56, 60], where qubits are formed using metastable states of neutral 171Yb atoms. In these atomic qubits single qubit, measurement, and idling errors are extremely unlikely [56, 72]; we hence neglect these for now. Instead, the dominant source of errors are two-qubit gate errors that arise during state preparation or in manipulations required to utilize the state as a resource. These errors primarily stem from a decay process that takes the system out of the computational Hilbert space and constitute experimentally detectable erasure errors. Since these decays only occur from one of the two computational states, re-initializing the system in this state can produce at most $Z$-errors.
@@
Here $e^{-}=|1\rangle\langle 0|$ and $e^{+}=|0\rangle\langle 1|$ are raising/lowering operators for the number $n^{e}_{j}$ of erasures on site $j$ (which lie outside the logical space), $Z_{j}$ is the Pauli-$Z$ operator in the logical space, and $\eta$ parametrizes the background erasure noise rate, with $\eta dt$ the error probability in a single application of the corresponding noise channel (dropping terms of order $(\eta dt)^{2}$). This channel has two crucial features: first, in this idealized limit, all errors are heralded; second, it has strong $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry.
@@
Here $\gamma$ is the correction rate, with a correction probability $\gamma dt$ in a single application of the correction channel $\epsilon_{C}(\rho)$.
Figure 2: Estimation of critical exponents ($L=512$). (a) Stabilizer defects: The running estimate of the exponent $\delta(t):=\log_{10}\frac{(0.5-n_{d})(t)}{(0.5-n_{d})(10t)}$ is plotted as a function of $1/t$. The order parameter at the critical point vanishes as $t^{-\delta_{d}}$ so $\lim_{t\rightarrow\infty}\delta_{d}(t,\eta=\eta_{c})=\delta_{d}$. In the absorbing (active) phase $\delta_{d}(t)\rightarrow\infty(0)$. The critical value of $\eta$ is estimated by identifying the curve that saturates to a constant value as $\frac{1}{t}\rightarrow 0$ and the corresponding constant value is the exponent $\delta_{d}$. The inset shows the scaling collapse for $(0.5-n_{d})t^{\delta}$ as a function of $t|\eta-\eta_{c}|^{\nu_{t}}$ when $\nu_{t}$ is tuned to $1.73$. (b) The string order exponents $\delta_{\Omega}$ and $\nu_{t}$ are estimated similarly. The data is obtained by averaging over $4\times 10^{4}$ Monte Carlo trajectories initialized in the cluster state and simulated upto $L^{1.5}$ MC sweeps (see SM [64] for details).
Figure 3: Finite non-erasure errors: (a) The $f_{e}-\eta$ phase diagram of the system. $\eta$ is the total error rate and $f_{e}$ is the fraction of errors that generate erasures. The color-plot shows the string-order of length $l=L/2$ evaluated at a finite time $t=2/\eta_{c}$. This half-chain string-order decays to zero exponentially with time for any $f_{e}<1$. The data is obtained by averaging over $5\times 10^{3}$ Monte Carlo runs. (b) The corresponding time constant $\tau_{\Omega}$ at $\eta=0.3<\eta_{c}$, obtained by a best fit to the numerical data, is plotted as a function of the non-erasure fraction $1-f_{e}$. The corresponding steady-state string-order is analyzed along the cut $\eta=0.3$ in (c). The characteristic length scale $\xi_{\Omega}$ is obtained by fitting the steady-state string-order of length $l$ to $f(l)\sim e^{-l/\xi_{\Omega}}$. The length scale $\xi_{\Omega}$ is plotted as a function of $1-f_{e}$ for various system sizes $L$.
@@
$d\rho/dt=\sum_{\alpha=\gamma,\eta}\sum_{s=1}^{4}\sum_{j=1}^{2L}L_{\alpha,s,j}\rho L^{\dagger}_{\alpha,s,j}+\frac{1}{2}\{L^{\dagger}_{\alpha,s,j}L_{\alpha,s,j},\rho\}$ (4)
@@
Stability.—We now consider stability of the SPT-ordered steady-state against local perturbations to its dynamics. Since string-order vanishes in the absence of strong symmetry even for a finite-depth local channel [49, 50], our SPT steady-state is unstable against symmetry-breaking $X$ errors. This is analogous to closed quantum systems, where there are no non-trivial bosonic phases in 1D in the absence of symmetry [62]. To study stability against strongly symmetric noise, we consider unheralded $Z$ errors which can arise experimentally from imperfect heralding or from unitary dynamics separating defects from the underlying erasures. We set the rate of heralded errors in Eq. (1) to $\eta f_{e}$, with unheralded errors described by the Lindblad operators $L_{f,1,j}=\sqrt{\eta(1-f_{e})}Z_{j}$. The total Pauli error rate is thus $\eta$, with $f_{e}$ the fraction of heralded errors. Fig. 3a shows the string-order at time $t=2/\eta_{c}$ as a function of the heralding fraction $f_{e}$ and the noise to correction ratio $\eta/\gamma$. At low noise rates and $f_{e}\sim 1$, the string-order at this time-scale is greatly enhanced relative to the unheralded model $(f_{e}=0)$. Fig. 3b shows that the string-order decays exponentially in time to its steady-state value with a characteristic time-scale $\tau_{\Omega}\sim 1/(1-f_{e})$. However, for $f_{e}<1$ our model does not realize a true steady-state SPT since the steady-state string-order decays exponentially with the string’s length (Fig. 3c), with a characteristic length scale $\xi_{\Omega}\sim 1/\sqrt{1-f_{e}}$. In the SM [64], we show that despite the lack of steady-state string-order, the absorbing state transition persists for $f_{e}<1$, and that in the phase where erasures have not proliferated, our correction protocol substantially reduces both $n_{d}$ and the maximum length of error strings in the steady-state. We also show that signatures of this critical point are detectable from the finite-time behavior of the string-order parameter, which exhibits a crossover from characteristic power-law decay to an exponential decay given by $\Omega(t)\sim t^{-\delta_{\Omega}}e^{-4\eta(1-f_{e})t}$ (see SM [64]).
@@
Discussion.—In this work, we have demonstrated that biased erasure noise permits the stabilization of nontrivial zero-temperature phases as steady-states under local Lindblad dynamics. Crucially, our experimentally motivated noise model can be exploited to effectively confine the errors which otherwise proliferate at any finite temperature or error rate (resulting in a trivial steady-state), thereby enabling local error correction into an SPT phase starting from generic initial mixed states. We have also shown that our protocol significantly enhances the lifetime of string-order in the presence of unheralded errors. Although we have focused on 1D SPT order here, we expect that our results extend more broadly to nontrivial ordered states which lack thermal stability, which we discuss in a forthcoming work.
@@
Acknowledgements: The authors are grateful to Tim Hsieh, Sanjay Moudgalya, Ramanjit Sohal, David Stephen, and Jeff Thompson for stimulating discussions. F.J.B. and S.C. are grateful for the support of NSF DMR 1928166. S.G. is supported in part by an Institute for Robust Quantum Simulation (RQS) seed grant. This material is based upon work supported by the Sivian Fund at the Institute for Advanced Study and the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0009988 (A.P.). The authors acknowledge the computational resources provided by the Minnesota Supercomputing Institute (MSI) at the University of Minnesota.
@@ I Review of the 1D $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ Cluster State SPT
Let us now study the system on open boundary conditions. In this case, the cluster entangler is given by $U=\prod_{j=1}^{L-1}CX_{j,j+1}$, such that
@@ I.1 Stability of SPT order under finite-depth local quantum channels
$\braket{\Omega_{ij}}_{\rho_{\mathcal{E}}}\sim(1-2p)^{i-j}\braket{\psi_{C}}{\Omega_{ij}}{\psi_{C}}\,,$ (S11)
@@ III Details of the Correction protocol
The error model also generates configurations of the form $\{1,1,0\}$ and $\{0,1,1\}$ where the intermediate site between an erasure and a non-erasure does not have a stabilizer defect. Such configurations result from the part of the erasure error channel that acts as $\mathbb{I}$ on the system. We can safely remove such isolated erasure records by applying transition operators $e^{+}_{j+1}$ or $e^{+}_{j-1}$. The corresponding Kraus operators are
@@
In the main text, we study the dynamics of a system initialized in the cluster state, which only generates a classical mixture of stabilizer eigenstates (refer to Sec. VII for further details). Note that if the measurement is not followed by a feedback operation, the eigenstates of the measurement operator remain unchanged. Anticipating this, we can drop the terms corresponding to $K_{0,1,0},K_{0,-1,0},K_{1,1,1}$ and $K_{1,-1,1}$ since they have no effect on the time evolution of the cluster state. Then, after taking the short time limit $dt\rightarrow 0$, we obtain the Lindblad master equation that represents the correction dynamics. The jump operators of the master equation encoded in Eqs. (2) and  (3) of the main text can be read off from the Kraus operators, which we reproduce here for the sake of completeness:

### v3

@@ Abstract
Figure 1: (a) Top: schematic showing stabilizers $X_{j-1}Z_{j}X_{j+1}$ (triangles) and locations of erasure errors (flags) relative to the qubits (circles). Bottom: the correction protocol for different configurations (the remainder of the system is unchanged). The left displays processes that move a stabilizer defect (filled triangle) past an adjacent erasure (filled flag); the right shows processes that shorten erasure strings which do not end on a stabilizer defect. (b) The steady-state expectation values of the string-order ($\Omega$), density of stabilizer defects ($n^{d}$), and the density of erasures ($n^{e}$) are shown as a function of the erasure noise rate $\eta/\gamma$ and clearly display a sharp transition. The data is shown for $L=128,256,512$ sites per sublattice (increasing system sizes from lighter to darker shades). The steady-state values are obtained by averaging over $5-10\times 10^{3}$ independent Monte Carlo trajectories initialized in the cluster state and simulated upto $2\times L^{1.5}$ MC sweeps.
@@
In this work, we show that for a particular class of experimentally relevant error models, strictly local error correction can stabilize one-dimensional symmetry protected topological (SPT) order in steady-state phases. These steady-state phases have a non-vanishing string-order parameter, ensuring that the mixed state remains a resource for quantum teleportation [37] and measurement-based quantum computation (MBQC) [38, 39]. In closed quantum systems, SPT order is stable against symmetric local perturbations, for which symmetry defects are locally created in pairs. In open quantum systems, only “strong” symmetries correspond to the conservation of symmetry charges within the system [40] and are hence required for defining mixed-state SPT order  [^1] Mixed-state order can be defined for $G\times H$ symmetric systems, where $G$ is a “weak” symmetry and $H$ is strong [41, 15, 42, 43, 44, 45, 46, 47, 48]. [49, 50, 51, 52, 41, 15, 42, 43, 53]. While pure SPT states are stable to strongly symmetric finite-depth local channels [49, 50, 54], here we demonstrate that mixed SPT phases can arise as stable steady-states under local Lindbladian evolution.
@@
Background on SPTs.—We will focus on stabilizing the 1D cluster state |$\ket{\psi}_{C}=\prod_{j=1}^{2L}CX_{j,j+1}\ket{\uparrow}^{\otimes 2L}$ (where $CX_{i,j}=\frac{1}{2}\left(1+X_{j}+X_{j+1}-X_{j}X_{j+1}\right)$ is a two-qubit conditional unitary), which is an SPT state protected by internal $G=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. |$\ket{\psi}_{C}$ is a representative stabilizer state within this SPT phase, satisfying $S_{j}\ket{\psi}_{C}=\ket{\psi}_{C}$ for all $j$, with stabilizers $S_{j}=X_{j-1}Z_{j}X_{j+1}$. While |$\ket{\psi}_{C}$ is short-range entangled, it cannot be connected to a $G$-symmetric paramagnetic ground state (i.e. a symmetric product state) through a $G$-symmetric finite-depth local unitary [62, 63, 64]. The resulting SPT order is encoded in the nonlocal “string-order parameters,” $\Omega^{a}_{i_{0},j_{0}}=X_{2i_{0}-1}\left(\prod_{i=i_{0}}^{j_{0}}Z_{2i}\right)X_{2j_{0}+1}$, $\Omega^{b}_{i_{0},j_{0}}=X_{2i_{0}}\left(\prod_{i=i_{0}}^{j_{0}}Z_{2i+1}\right)X_{2j_{0}+2}$, with global symmetry generators $\mathbb{Z}_{2}^{(a)}=\prod_{i=1}^{L}Z_{2i}$, $\mathbb{Z}_{2}^{(b)}=\prod_{i=1}^{L}Z_{2i-1}$ for a system with $L$ unit cells. At generic points within the phase, when the interval $(i_{0},j_{0})$ spans the entire system, and thus has a definite symmetry charge, the string operators reduce to long-range two-point correlation functions of operators localized at the boundaries, resulting in edge states with fractionalized symmetry (see Supplemental Material (SM) [65] for details). This leads to entanglement features that are a resource for teleportation or MBQC throughout the SPT phase [66, 67, 68, 69, 70, 71, 37, 72].
Noise Model.— Heralded noise occurs in systems where errors leave a signature that can be detected without disturbing wave-functions in the logical space. Here, we consider a particular model of heralded noise, namely biased erasure noise, which naturally exhibits strong $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. This model is specifically motivated by the Rydberg atom setup considered in Ref. [57, 61], where qubits are formed using metastable states of neutral 171Yb atoms. In these atomic qubits single qubit, measurement, and idling errors are extremely unlikely [57, 73]; we hence neglect these for now. Instead, the dominant source of errors are two-qubit gate errors that arise during state preparation or in manipulations required to utilize the state as a resource. These errors primarily stem from a decay process that takes the system out of the computational Hilbert space and constitute experimentally detectable erasure errors. These erasures can be detected without affecting the logical state of the qubit, yielding a model with heralded noise. Since these decays only occur from one of the two computational states, re-initializing the system in this state can produce at most $Z$-errors.
@@
Here $e^{-}=|1\rangle\langle 0|$ and $e^{+}=|0\rangle\langle 1|$ are raising and lowering operators for the number $n^{e}_{j}$ of erasures on site $j$ (which lie outside the logical space), $Z_{j}$ is the Pauli-$Z$ operator in the logical space, and $\eta$ parametrizes the background erasure noise rate, with $\eta dt$ the error probability in a single application of the corresponding noise channel (dropping terms of order $(\eta dt)^{2}$). This channel has two crucial features: first, in this idealized limit, all errors are heralded; second, it has strong $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry.
@@
Here $\gamma$ is the correction rate, with a correction probability $\gamma dt$ in a single application of the correction channel $\epsilon_{C}(\rho)$. A circuit implementing these correction operators is given in [65].
Figure 2: Estimation of critical exponents ($L=512$). (a) Stabilizer defects: The running estimate of the exponent $\delta(t):=\log_{10}\frac{(0.5-n^{d}(t))}{(0.5-n^{d}(10t))}$ is plotted as a function of $1/t$. The order parameter at the critical point vanishes as $t^{-\delta_{d}}$ so $\lim_{t\rightarrow\infty}\delta_{d}(t,\eta=\eta_{c})=\delta_{d}$. In the absorbing (active) phase $\delta_{d}(t)\rightarrow\infty(0)$. The critical value of $\eta$ is estimated by identifying the curve that saturates to a constant value as $\frac{1}{t}\rightarrow 0$ and the corresponding constant value is the exponent $\delta_{d}$. The inset shows the scaling collapse for $(0.5-n^{d})t^{\delta}$ as a function of $t|\eta-\eta_{c}|^{\nu_{t}}$ after initial transient when $\nu_{t}$ is tuned to $1.73$. (b) The string order exponents $\delta_{\Omega}$ and $\nu_{t}$ are estimated similarly. The data is obtained by averaging over $4\times 10^{4}$ Monte Carlo trajectories initialized in the cluster state and simulated upto $L^{1.5}$ MC sweeps (see SM [65] for details).
Figure 3: Finite non-erasure errors: (a) The $f_{e}-\eta$ phase diagram of the system. $\eta$ is the total error rate and $f_{e}$ is the fraction of errors that generate erasures. The color-plot shows the string-order of length $l=L/2$ evaluated at a finite time $t=2/\eta_{c}$, which is chosen to maximize the contrast between the active phase (where string order remains strong at this time) and the absorbing phase (where it has all but vanished at time $2/\eta_{c}$). This half-chain string-order decays to zero exponentially with time for any $f_{e}<1$ and vanishes in the steady-state. The data is obtained by averaging over $5\times 10^{3}$ Monte Carlo runs. (b) The corresponding time constant $\tau_{\Omega}$ at $\eta=0.3<\eta_{c}$, obtained by a best fit to the numerical data (see [65] for details), is plotted as a function of the non-erasure fraction $1-f_{e}$. The corresponding steady-state string-order is analyzed along the cut $\eta=0.3$ in (c). The characteristic length scale $\xi_{\Omega}$ is obtained by fitting the steady-state string-order of length $l$ to $f(l)\sim e^{-l/\xi_{\Omega}}$. The length scale $\xi_{\Omega}$ is plotted as a function of $1-f_{e}$ for various system sizes $L$.
@@
$d\rho/dt=\sum_{\alpha=\gamma,\eta}\sum_{s=1}^{4}\sum_{j=1}^{2L}L_{\alpha,s,j}\rho L^{\dagger}_{\alpha,s,j}-\frac{1}{2}\{L^{\dagger}_{\alpha,s,j}L_{\alpha,s,j},\rho\}$ (4)
@@
Stability.—We now consider stability of the SPT-ordered steady-state against local perturbations to its dynamics. Since string-order vanishes in the absence of strong symmetry even for a finite-depth local channel [49, 50], our SPT steady-state is unstable against symmetry-breaking $X$ errors. This is analogous to closed quantum systems, where there are no non-trivial bosonic phases in 1D in the absence of symmetry [63]. To study stability against strongly symmetric noise, we consider unheralded $Z$ errors which can arise experimentally from imperfect heralding or from unitary dynamics separating defects from the underlying erasures. We set the rate of heralded errors in Eq. (1) to $\eta f_{e}$, with unheralded errors described by the Lindblad operators $L_{f,1,j}=\sqrt{\eta(1-f_{e})}Z_{j}$. The total Pauli error rate is thus $\eta$, with $f_{e}$ the fraction of heralded errors. Fig. 3a shows the string-order at time $t=2/\eta_{c}$ as a function of the heralding fraction $f_{e}$ and the noise to correction ratio $\eta/\gamma$. At low noise rates and $f_{e}\sim 1$, the string-order at this time-scale is greatly enhanced relative to the unheralded model $(f_{e}=0)$. Fig. 3b shows that the string-order decays exponentially in time to its steady-state value with a characteristic time-scale $\tau_{\Omega}\sim 1/(1-f_{e})$. However, for $f_{e}<1$ our model does not realize a true steady-state SPT since the steady-state string-order decays exponentially with the string’s length (Fig. 3c), with a characteristic length scale $\xi_{\Omega}\sim 1/\sqrt{1-f_{e}}$. In the SM [65], we show that despite the lack of steady-state string-order, the absorbing state transition persists for $f_{e}<1$, and that in the phase where erasures have not proliferated, our correction protocol substantially reduces both $n^{d}$ and the maximum length of error strings in the steady-state. We also show that signatures of this critical point are detectable from the finite-time behavior of the string-order parameter, which exhibits a crossover from characteristic power-law decay to an exponential decay given by $\Omega(t)\sim t^{-\delta_{\Omega}}e^{-4\eta(1-f_{e})t}$ (see SM [65]).
@@
Discussion.—In this work, we have demonstrated that biased erasure noise permits the stabilization of nontrivial zero-temperature phases as steady-states under local Lindblad dynamics. Crucially, our experimentally motivated noise model can be exploited to effectively confine the errors which otherwise proliferate at any finite temperature or error rate (resulting in a trivial steady-state), thereby enabling local error correction into an SPT phase starting from generic initial mixed states. We have also shown that our protocol significantly enhances the lifetime of string-order in the presence of unheralded errors. Although we have focused on 1D SPT order here, our results extend more broadly to nontrivial ordered states in higher dimensions which lack thermal stability; in particular, heralding enables local error-correction into 2D topologically ordered steady-state phases, which we discuss in Ref. [80].
@@
Acknowledgements: The authors are grateful to Tim Hsieh, Sanjay Moudgalya, Ramanjit Sohal, David Stephen, and Jeff Thompson for stimulating discussions. F.J.B. and S.C. are grateful for the support of NSF DMR-1928166 and NSF DMR-2313858. S.G. is supported in part by an Institute for Robust Quantum Simulation (RQS) seed grant. This material is based upon work supported by the Sivian Fund and the Paul Dirac Fund at the Institute for Advanced Study and the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0009988 (A.P.). The authors acknowledge the computational resources provided by the Minnesota Supercomputing Institute (MSI) at the University of Minnesota.
@@ I Review of the 1D $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ Cluster State SPT
Let us now study the system on open boundary conditions. In this case, the cluster entangler is given by $U=\prod_{j=1}^{2L-1}CX_{j,j+1}$, such that
@@ I.1 Stability of SPT order under finite-depth local quantum channels
$\braket{\Omega_{ij}}_{\rho_{\mathcal{E}}}\sim(1-2p)^{|i-j|}\braket{\psi_{C}}{\Omega_{ij}}{\psi_{C}}\,,$ (S11)
@@ III Details of the Correction protocol
Figure S1: The quantum circuit that implements the correction protocol using the native gate-set (Hadamard ($H$) and controlled-Z) of the Rydberg atom platform. The ancilla qubit, initialized in the $|0\rangle_{\mathrm{ancilla}}$ state, is sequentially entangled with system qubits located at $j-1,j$, and $j+1$. A measurement of the ancilla in the computational basis performs a measurement of the stabilizer $S_{j}=X_{j-1}Z_{j}X_{j+1}$ on the system qubits. The feedback $Z_{j-1}$ is applied conditioned on the measurement outcome $m\in\{0,1\}$. Subsequently, the flag at site $j-1$ is lowered from $1$ to 00 at the end of the process. This particular circuit shows the application of the jump operator $\frac{1}{2}e^{+}_{j-1}Z^{m}_{j-1}(1+(-1)^{m}S_{j})(1-n^{e}_{j+1})$ when the defect is corrected in the direction of $j-1$. The rest of the correction operators can be designed analogously. Note that while the total measurement+feedback operator obeys the strong symmetry, the component gate $H$ is not strongly symmetric and can hence potentially convert $Z$ errors into $X$ errors. It would be desirable to find ways to develop native gates that obey the symmetry condition and preserve the noise bias. The circuit diagram is drawn using the Quantikz pacakge [85].
@@
The quantum circuit that implements this measurement and feedback operation is shown in Fig. (S1). The error model also generates configurations of the form $\{1,1,0\}$ and $\{0,1,1\}$ where the intermediate site between an erasure and a non-erasure does not have a stabilizer defect. Such configurations result from the part of the erasure error channel that acts as $\mathbb{I}$ on the system. We can safely remove such isolated erasure records by applying transition operators $e^{+}_{j+1}$ or $e^{+}_{j-1}$. The corresponding Kraus operators are
@@
In the main text, we study the dynamics of a system initialized in the cluster state, which only generates a classical mixture of stabilizer eigenstates (refer to Sec. IV for further details). Note that if the measurement is not followed by a feedback operation, the eigenstates of the measurement operator remain unchanged. Anticipating this, we can drop the terms corresponding to $K_{0,1,0},K_{0,-1,0},K_{1,1,1}$ and $K_{1,-1,1}$ since they have no effect on the time evolution of the cluster state. Then, after taking the short time limit $dt\rightarrow 0$, we obtain the Lindblad master equation that represents the correction dynamics. The jump operators of the master equation encoded in Eqs. (2) and  (3) of the main text can be read off from the Kraus operators, which we reproduce here for the sake of completeness: